[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 11, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 11 13:02:34 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. Erratic mouse cursor motion issues on the Cumulus, Shiner and Ballyhoo  
N-AWIPS workstations. This seems to be caused by a mismatch between the  
graphics card and the version of the Nvidia driver software installed,  
which causes the video RAM to run low. TSB is investigating whether a more  
stable version of the video driver can be installed on these systems.

2. For three nights in a row (Saturday-Monday), the product tracker failed  
to notice that the 03Z issuance of the OFFNT3 (offshore waters forecast):
http://www.nhc.noaa.gov/text/MIAOFFNT3.shtml
had been sent out. It did work correctly for today's 03Z OFFNT3. TSB will  
monitor the situation.
HSU:
1. Senior Hurricane Specialist Stacy Stewart did not receive an automatic  
comeback copy (print) of the TCPAT5 product (Atlantic public advisory 29)  
for TS Nicole, even though it posted to AWIPS, NWX, and the NHC web page:
http://www.nhc.noaa.gov/archive/2016/al15/al152016.public.029.shtml?
TSB will investigate.
Other items:
1. NHC has requested and received approval to extend GOES-East Rapid  
Scanning Operations until 1514 UTC (11:14 EDT) on Wednesday, 10/12/16 for  
the purpose of tracking and monitoring TS Nicole.

NHC RFCs:
RFC #204 – rsync_info.pl v1.0.1 - Add Cray IPs and ATCF database paths to  
the remote host hash to begin syncing the ATCF database up to the Crays for  
future development on the new systems. Directory paths have already been  
set up. To be implemented on Tuesday, October 11 at 1230Z (8:30 am EDT).

RFC #2422 - Modify the LDM tables to route one more aircraft WMO header to  
disk on WCOSS. This is in support of NHC's mission. To be implemented on  
Tuesday, October 11 at 1430Z (10:30 am EDT).

RFC #753 (OPC) - Update the ETSS surge web images to be hourly and include  
a choice for total water level at MDL's request. Update both ETSS and  
ESTOFS images to include hourly 0.25 degree GFS EMSL data, replacing 3  
hourly 0.5 degree PMSL data. Update the web page to display the new images  
and process run time monitoring for the new scripts. To be implemented on  
Tuesday, October 1 at 1500Z (11 am EDT).


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161011/8cedaa1c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list